# Do we really have to filter out random noise in pre-training data for language models?

### v1

@@ D.1 Pre-training Dataset
Figure 8: The prior distribution of tokens in the data from (a) OpenWebText, (b) random noise, and (c) Gaussian noise.
OpenWebText Dataset. The OpenWebText dataset (Gokaslan et al., 2019) is a large-scale corpus of English text data, developed to serve as an open-access alternative to proprietary dataset WebText that is utilized by OpenAI for training their GPT-2 models. This dataset originates from the analysis of outbound links clicked on Reddit, undergoing multiple stages of filtering to exclude non-English content, duplicate entries, copyrighted materials, and texts lacking in quality. These links generally direct to web pages available to the public, often shared or debated on Reddit, thereby covering a broad spectrum of subjects that mirror online popular interests and discussions. The dataset includes roughly 18 million documents, amounting to about 20GB of compressed plain text data in uint16 format. Since measures have been implemented to ensure the dataset’s reliability by filtering out unsuitable content, we consider it a clean and noise-free dataset. Figure 8(a) illustrates the distribution of internal tokens.
Random Noise. To simulate the distribution of random gibberish that crawlers might retrieve from the Internet due to various reasons, we manually searched and collected a few websites containing such gibberish and also opened normally functioning websites using different decoding methods to observe the distribution of tokens. We found that, while the distribution of tokens appeared disorganized, their prior probabilities were not evenly distributed. Instead, several tokens had notably high probabilities, which is similar to that observed in the clean data as shown in Figure 8(a). Thus, on the basis of uniformly distributed random noise, we increased the frequency of certain tokens and then randomized them again. The resulting distribution is illustrated in Figure 8(b). It can be seen that while maintaining an overall uniform distribution, the frequency of tokens with IDs ranging from 0 to 1000 is higher which closely mirrors real-world scenarios.
Gaussian Noise. Given the diversity and unpredictability of real-world data distributions, we also artificially generated random noise with a prior probability that follows a Gaussian distribution, as shown in Figure 8(c). The rationale behind choosing the Gaussian distribution is that the noise in many real-world systems can be approximated by it. Additionally, we set the standard deviation $\sigma=500$ to simulate scenarios where random noise exhibits sharp peaks.
@@ D.2 Training Details of GPT-2
Our work is based on the source code of nanoGPT [^4] https://github.com/karpathy/nanoGPT. Specifically, we utilized the GPT-2 tokenizer with vocabulary size $V=50256$ to process the OpenWebText dataset, and then appended randomly generated noise to the end of the training set before commencing training. The model’s context length is set to $L=1024$, with an embedding dimension $d=768$. The GPT-2 model consists of 12 self-attention layers, totaling approximately 124 million parameters. For optimization, we employed AdamW (Loshchilov, 2017; Xie et al., 2022) with a learning rate of 6e-4, weight decay of 0.1, and $\beta$ values of 0.9 and 0.95 for $\beta_{1}$ and $\beta_{2}$, respectively. A cosine annealing scheduler was used to gradually adjust the learning rate down to 6e-5. We configured the batch size to 16, with a gradient accumulation step of 40, allowing each iteration to process 655,360 tokens (16 * 40 * 1024). Training proceeded for a total of 300,000 iterations.
D.3 Synthetic Results about Multilingual Models
@@
D.4 Hardware
We conducted the pre-training process on a server equipped with 8 NVIDIA GeForce RTX 4090 GPUs. It takes approximately 70 hours to train one model using eight 4090 GPUs, so pre-training five GPT-2 models in total requires 2,800 GPU hours.
D.5 Additional Results on GPT-2 774M and BERT-base
To further validate the impact of random noise, we scaled up the parameters of GPT-2 from 124M to 774M and conducted pre-training on a dataset containing 5% random noise. The results are presented in Table 8.
Noise Percentage 10,000 iters 50,000 iters 100,000 iters
0% 3.04 2.89 2.82
5% 3.06 2.92 2.83
Table 8: NTP loss during GPT-2 774M training
These findings sufficiently demonstrate that our conclusions can be extended to models with larger scales, as our theoretical framework does not hinge upon model performance metrics.
Additionally, experiments in Section 4 involve BERT and vision models to assess the generalizability of the proposed local gradient matching function. This investigation is somewhat tangential to our primary focus—the impact of random noise on training language models using autoregressive approaches. Nonetheless, we pre-trained the BERT-base model under 5% noise conditions. The outcomes are illustrated below.
Noise Percentage 1M iters 5m iters 10m iters
0% 3.21 2.94 2.76
5% 3.23 2.97 2.79
Table 9: NTP loss during BERT-base training.
These results align well with our theoretical predictions.
For vision models, pre-training outcomes can be found in Chen et al. (2024). Specifically, the authors simulated label noise by randomly replacing ImageNet sample labels and tested on 14 downstream tasks. The slight decrease in accuracy further corroborates our hypothesis across different types of models and datasets, underscoring the robustness of our approach against various levels of noise.
@@ E.1.1 Datasets
SST-2 (Socher et al., 2013) 2 6.92k 1.82k
SST-fine (Chen and Manning, 2014) 5 8.54k 2.21k
20newsgroup (Zhang et al., 2019) 20 11.3k 7.53k
CR (Hu and Liu, 2004) 2 3.39k 376
BBC (Samuels and Mcgonical, 2020) 5 1.23k 1k
Balanced COPA (Kavumba et al., 2020) 2 1k 500
MRPC (Dolan and Brockett, 2005) 2 3.67k 1.73k
WiC (Pilehvar and Camacho-Collados, 2019) 2 5.43k 1.4k
@@ E.1.2 Prompts
Since classification tasks can be processed as seq2seq tasks by adding prompts (Sutskever, 2014; Saunshi et al., 2021), we design a unique prompt for each dataset and task. This approach transforms the inputs into a format that large language models can process. The specific designs are shown in Table 11.
@@ E.1.3 Hyperparameters
For all experiments shown in Table 1, we set $\gamma$ in Equation (3) to be 0.01 and $\lambda$ in Equation (6) to be 0.15. Following the setup as described by Saunshi et al. (2021), for each dataset, we conduct a grid search on the validation set to identify the optimal learning rate and batch size. We train for a total of ten epochs with the learning rate ranging within {3e-4, 6e-4} and batch size options including {8, 12, 16, 32}. For samples without a designated validation set, we randomly select 10% of the training set samples to form a validation set for the purpose of selecting the best parameters.
@@ E.2.1 Datasets
We select 14 image classification datasets, which serve as a common benchmark for evaluating model performance in the vision community (Zhou et al., 2022b; Chen et al., 2024). Specific information about these 14 datasets is provided in Table 12.
@@
StanfordCars (Krause et al., 2013) 196 8144 8041
Caltech101 (Fei-Fei et al., 2004a) 102 3060 6084
CIFAR-10 (Krizhevsky et al., 2009) 10 50000 10000
CIFAR-100 (Krizhevsky et al., 2009) 100 50000 10000
DTD (Cimpoi et al., 2014) 47 1880 1880
EuroSAT (Helber et al., 2019) 10 21600 5400
FGVCAircraft (Maji et al., 2013) 102 6667 $3$
Flowers102 (Nilsback and Zisserman, 2008) 102 2040 6149
Food101 (Fei-Fei et al., 2004b) 101 75750 25250
OxfordPet (Parkhi et al., 2012) 37 3680 3669
PatchCamelyon (Veeling et al., 2018) 2 262144 32768
RESISC45 (Cheng et al., 2017) 45 25200 6300
Rendered SST2 (Socher et al., 2013) 2 6920 1821
SVHN (Netzer et al., 2011) 10 73257 26032
@@ E.2.2 Models
ResNetv2-152x2 (He et al., 2016) ImageNet-21K (Ridnik et al., 2021) 321.7M
Swin-L (Liu et al., 2021) ImageNet-21K 196.7M
ConvNext-L (Woo et al., 2023)
Laion-2B (Schuhmann et al., 2022)
@@ E.2.3 Hyperparameters
In our study, similar to the approach detailed in Chen et al. (2024), we primarily contrast our proposed method with MLP and LP tuning. For the optimization process, we employ AdamW for fine-tuning the modules over 30 epochs, utilizing a cosine learning rate scheduler. Specifically, for LP, we configure the learning rate at 0.1 without applying any weight decay. In contrast, both the MLP tuning and our method use a more conservative learning rate of 0.001 alongside a weight decay of 1e-4.

### v2

@@ D.1 Pre-training Dataset
OpenWebText Dataset. The OpenWebText dataset [13] is a large-scale corpus of English text data, developed to serve as an open-access alternative to proprietary dataset WebText that is utilized by OpenAI for training their GPT-2 models. This dataset originates from the analysis of outbound links clicked on Reddit, undergoing multiple stages of filtering to exclude non-English content, duplicate entries, copyrighted materials, and texts lacking in quality. These links generally direct to web pages available to the public, often shared or debated on Reddit, thereby covering a broad spectrum of subjects that mirror online popular interests and discussions. The dataset includes roughly 18 million documents, amounting to about 20GB of compressed plain text data in uint16 format. Since measures have been implemented to ensure the dataset’s reliability by filtering out unsuitable content, we consider it a clean and noise-free dataset.
@@ D.2 Training Details of GPT-2
Our work is based on the source code of nanoGPT [^3] https://github.com/karpathy/nanoGPT. Specifically, we utilized the GPT-2 tokenizer with vocabulary size $V=50256$ to process the OpenWebText dataset, and then appended randomly generated noise to the end of the training set before commencing training. The model’s context length is set to $L=1024$, with an embedding dimension $d=768$. The GPT-2 model consists of 12 self-attention layers, totaling approximately 124 million parameters. For optimization, we employed AdamW [106, 107] with a learning rate of 6e-4, weight decay of 0.1, and $\beta$ values of 0.9 and 0.95 for $\beta_{1}$ and $\beta_{2}$, respectively. A cosine annealing scheduler was used to gradually adjust the learning rate down to 6e-5. We configured the batch size to 16, with a gradient accumulation step of 40, allowing each iteration to process 655,360 tokens (16 * 40 * 1024). Training proceeded for a total of 300,000 iterations.
D.3 Experiments on Other Text Corpus
To further investigate the impact of random noise on model generalization, we evaluate the next-token prediction loss of the trained models on data crawled from arXiv and Wikipedia. The results are illustrated in Figure 7. Surprisingly, models trained with added noise outperformed those trained on $P^{c}$. This counterintuitive finding aligns with previous work in visual domains [85], suggesting that incorporating random noise into training sets might enhance model robustness. Additionally, we observe that the performance of models subjected to Gaussian noise varies across different datasets. These observations warrant further investigation.
D.4 Synthetic Results about Multilingual Models
@@
D.5 Hardware
We conducted the pre-training process on a server equipped with 8 NVIDIA GeForce RTX 4090 GPUs. It takes approximately 70 hours to train one model using eight 4090 GPUs, so pre-training five GPT-2 models in total requires 2,800 GPU hours. Pre-training a 2.7B model takes about a week on 4 NVIDIA A100 GPUs.
@@ E.1.1 Datasets
SST-2 [108] 2 6.92k 1.82k
SST-fine [109] 5 8.54k 2.21k
20newsgroup [110] 20 11.3k 7.53k
CR [111] 2 3.39k 376
BBC [112] 5 1.23k 1k
Balanced COPA [113] 2 1k 500
MRPC [114] 2 3.67k 1.73k
WiC [115] 2 5.43k 1.4k
@@ E.1.2 Prompts
Since classification tasks can be processed as seq2seq tasks by adding prompts [116, 24], we design a unique prompt for each dataset and task. This approach transforms the inputs into a format that large language models can process. The specific designs are shown in Table 9.
@@ E.1.3 Hyperparameters
For all experiments shown in Table 1, we set $\gamma$ in Equation (5) to be 0.01 and $\lambda$ in Equation (8) to be 0.15. Following the setup as described by [24], for each dataset, we conduct a grid search on the validation set to identify the optimal learning rate and batch size. We train for a total of ten epochs with the learning rate ranging within {3e-4, 6e-4} and batch size options including {8, 12, 16, 32}. For samples without a designated validation set, we randomly select 10% of the training set samples to form a validation set for the purpose of selecting the best parameters.
@@ E.2.1 Datasets
We select 14 image classification datasets, which serve as a common benchmark for evaluating model performance in the vision community [97, 17]. Specific information about these 14 datasets is provided in Table 10.
@@
StanfordCars [117] 196 8144 8041
Caltech101 [118] 102 3060 6084
CIFAR-10 [119] 10 50000 10000
CIFAR-100 [119] 100 50000 10000
DTD [120] 47 1880 1880
EuroSAT [121] 10 21600 5400
FGVCAircraft [122] 102 6667 $3$
Flowers102 [123] 102 2040 6149
Food101 [124] 101 75750 25250
OxfordPet [125] 37 3680 3669
PatchCamelyon [126] 2 262144 32768
RESISC45 [127] 45 25200 6300
Rendered SST2 [108] 2 6920 1821
SVHN [128] 10 73257 26032
@@ E.2.2 Models
ResNetv2-152x2 [132] ImageNet-21K [133] 321.7M
Swin-L [134] ImageNet-21K 196.7M
ConvNext-L [135]
Laion-2B [136]
@@ E.2.3 Hyperparameters
In our study, similar to the approach detailed in [17], we primarily contrast our proposed method with MLP and LP tuning. For the optimization process, we employ AdamW for fine-tuning the modules over 30 epochs, utilizing a cosine learning rate scheduler. Specifically, for LP, we configure the learning rate at 0.1 without applying any weight decay. In contrast, both the MLP tuning and our method use a more conservative learning rate of 0.001 alongside a weight decay of 1e-4.
@@ E.3 Runtime Analysis
Appendix F Limitations
Firstly, due to limitations in computational resources and costs, we pre-train only the GPT-2 124M and 2.7B model on the OpenWebText dataset. Additionally, the types of noise considered are limited to uniform and Gaussian distributions. However, based on Proposition 1, we argue that training GPT-2 on the Synthetic OpenWebText dataset is sufficient to uncover the essence of the issue, as Proposition 1 makes no assumptions about data distribution or model architecture.